# Neutral residues: Revisiting adapters for model extension

### v1

@@ Abstract
We address the problem of extending a pretrained large language model to a new domain that was not seen at training time, like adding a language for which the original model has seen no or little training data. Popular solutions like fine-tuning or low-rank adaptation are successful at domain adaptation, but formally they do not add any extra capacity and degrade the performance in the original domain.
Our paper analyzes this extension problem under three angles: data, architecture and training procedure, which are advantageously considered jointly. In particular, we improve adapters and make it possible to learn an entire new language while ensuring that the output of the neural network is almost unchanged in the original domain. For this purpose, we modify the new residual blocks in a way that leads each new residual block to output near-zeros in the original domain.
This solution of neutral residues, which borrows architectural components from mixture of experts, is effective: with only 20% extra learnable weights compared to an original model trained on English, we get results that are significantly better than concurrent approaches (fine-tuning, low-rank or vanilla adapters) in terms of the trade-off between learning a new language and not forgetting English.
Figure 1: Learning or Forgetting? Fine-tuning a model reduces the model performance on the original task. Potential trade-offs are governed by the learning rate, with extreme cases being (top-left) the original backbone and (bottom-right) finetuning on the new data with a large learning rate.
LoRA or vanilla adapters mitigate catastrophic forgetting but only to some extent. Our method significantly improves this compromise. Detailed results in Section 4.
@@ 1 Introduction
The dominating strategy for producing foundation models involves training from scratch on a large collection of data, typically trillions of tokens, covering multiple domains. Training such models is extremely costly in terms of computing resources. In view of the explosion of the number of gigantic models produced in the last years (the HuggingFace model repository claims to host about one million models), this training paradigm raises the problem of the economical and ecological sustainability of the model production pipelines. As an example, training the Llama 3 model is estimated by multiple sources to cost hundreds of millions dollars, mostly due to the training on 24,000 flagship GPUs H100 for months111 https://www.theinformation.com/articles/ten-gifts-nvidia-gave-its-investors , not counting the hidden costs of exploration.
@@
In contrast, the adapters introduced by Rebuffi et al. (2017) extend a pretrained model by adding new parameters. When transferring to a new domain or task, only these new weights are trained while those of the original backbone are frozen. This extension strategy, which was initially introduced in computer vision with convolutional neural networks, was subsequently adapted to textual transformers by Houlsby et al. (2019). Adapters benefit from the initial pretrained model and hence exhibit competitive transfer performance. They enjoy several additional properties: Only a subset of weights is trained. Most importantly, they formally add some capacity (unlike linear additive adapters like LoRA) and hence do not suffer from the aforementioned limitations. However, adapters still suffer significant forgetting in their current form and are therefore not sufficient for our problem.
In our paper, we build upon adapters and, accordingly, increase the capacity of the model by adding feed-forward blocks to the model, while offering a modular design (Pfeiffer et al., 2020). In particular, we consider the use-case where we add a new language to a pretrained model. We measure how the extended model performs both on the training criterion, namely perplexity, but also on downstream tasks such as question answering.
We then analyze how to optimize the compromise between learning a new language, and not forgetting the initial knowledge of the pretrained model. We study the impact of (1) training data; (2) training strategy, such as losses specifically intended at reducing the forgetting, and (3) architecture. This leads us to point out several important factors for successfully extending a model with adapters:
  • Data: When learning adapters, a simple way to drastically reduce the forgetting is to keep training with a small fraction of data similar to the original distribution.
  • Architecture: An adapter gating mechanism is an effective way to ensure that the network distinguishes when it should operate as the original neural network or when the new blocks are desirable to process the input data.
@@
  • Training: The gating mechanism is advantageously informed in a supervised manner. We proposed two strategies in that respect, both introducing a local loss: The first one is inspired by Mixture-of-Experts and involves an explicit domain classifier at each block. The second one involves a sparsity loss whose objective is to ensure that the residual connections output near-zero values when the input follows the pretraining distribution.
Overall, we show that multiple ingredients are advantageously intertwined to ensure that the extended model still approximately implements the original function when the input is from the original distribution, without being too constrained to incorporate a large amount of new knowledge. In our experiments, these conclusions are validated with three models by extending the pretrained model with two languages.
@@ 2 Related Work
Finetuning is however not sufficient when a large body of knowledge must be incorporated to the network. In this section we review a body of literature related to this problem and to our work. This includes solutions that add new knowledge into a pretrained network, as well as methods that incorporate gating mechanisms such as mixtures of experts (MoE).
@@
Enhancing language model capabilities using module based training has gained interest in recent years due to the high cost of full finetuning. Those methods are known as parameter efficient finetuning methods (PEFT; Houlsby et al. (2019); Hu et al. (2022); Li & Liang (2021)). Unlike full finetuning, they only require a limited amount of memory. Houlsby et al. (2019) proposed to insert small bottleneck layers, known as Adapters, within each transformer layer, allowing models to be finetuned by training only a small fraction of the parameters, while keeping the original model frozen. There is a large body a literature on adapters, see the overview by Pfeiffer et al. (2023). Low-Rank Adaptation (LoRA) by Hu et al. (2022) adds trainable low-rank matrices to transformer layers while keeping the original weights frozen.
@@
Several studies have investigated more advanced methods to address forgetting in continual learning (Biderman et al. (2024); Wu et al. (2024); Li et al. (2024); Zhong et al. (2024); Riemer et al. (2019); Li & Hoiem (2017)). Biderman et al. (2024) show that using LoRA helps reducing forgetting but is less efficient in learning than full finetuning especially for distributions far from the LLM pretraining distribution. Wu et al. (2024) proposed a continual learning method for LLMs, which amounts to adding new intermediate transformer blocks in an interleaved manner. This enables the injection of new knowledge while preserving the initial capabilities.
Mixture-of-Experts.
@@
MoEs have a long history in neural networks (Yuksel et al., 2012), for instance with set of classifiers where each individual classifier was specialized in a different task or aspects of data. Residual architectures (He et al., 2016) and in particular transformers (Vaswani et al., 2017) have renewed the interest in this strategy. In this context, they amount to replacing the standard feed-forward networks (FFN) of the transformer blocks by a collection of FFNs, referred to as experts. A gating mechanism selectively activates a subset of relevant experts for a given input. This technique has been widely used to reduce the computational cost of inference in large language models (Jiang et al., 2024; Xue et al., 2024) with a focus on the design of an expert-balancing loss to promote equitable token distribution across experts (Shazeer et al., 2017; Pfeiffer et al., 2023).
@@
Recent works have explored utilizing MoE’s gating mechanism to mitigate catastrophic forgetting during the continual learning phase of language models. LorRAMoE (Dou et al., 2024) employs LoRAs as experts, integrating them with a gating mechanism to improve performance on specific downstream tasks through Supervised finetuning (SFT). LoRAMoE mitigates catastrophic forgetting in the case of SFT, but to our knowledge its effectiveness has not been studied in a broader context. In contrast, two recent parallel concurrent works are closer to our proposal: Zhong et al. (2024) augment Mixture-of-Experts LLMs with new modalities by addition of new experts to a pretrained mixture of experts. Similarly, Li et al. (2024) add an MoE layer in parallel to the FFN of the transformer block to learn multilingual capabilities and mitigate catastrophic forgetting during Continual Learning. Our solution shares some similarities with this work by employing mixed data training. Yet Li et al. (2024) do not integrate an explicit sparsity criterion, and need to weight differently the blocks of the pretrained model with those of the gated adapters. They also consider multiple elements in the mixture and apply a softmax for the rooting mechanism. Since at least one expert is activated in their mixture, this selection mechanism also prevents the capacity of the model to produce 0-valued outputs. Our solution of neutral residues with local loss is more effective than simply gating adapters, as shown in our experimental Section 4.
@@ 3 Neutral Residues with gated adapters
The first adapters by Rebuffi et al. (2017) were initially introduced in the context of convolutional residual networks. Their motivation was to adapt a backbone to multiple tasks by finetuning on a limited amount of data, rather than adding a large body of knowledge. Therefore they consider a limited number of trainable parameters, typically less than 10% additional weights compared to the original backbone. In the context of language models, Houlsby et al. (2019) add residual feed-forward block sequentially after each multi-head attention and the original FFN. In our case and as shown in Figure 2, we add more capacity (typically 20%) and adopt parallel adapters as we do not observe a noticeable difference compared to serial adapters. This concurs with observations by Touvron et al. (2022) that blocks can be parallelized pairwise if vision transformers are large enough.
@@
Figure 2: Architectural design adopted for neutral residues: (left) We add the adapter block in parallel to the FFN. (right) Our gated adapter block includes a gate with a ELU linearity (see Section 4.4 for a comparison with a sigmoidal gating). At training time, the trainable parameters are represented by red blocks. The weights of the blue blocks are frozen. The output of the adapter is optimized with a local loss such that the output is sparse if the input follows the pretraining distribution.
Near-identical initialization.
@@
FFN vs multi-head attention (MHA).
In our experiments carried out for adding a new language, it is more effective to add extra weights in the form of FFN than in MHA blocks, for a given number of additional parameters. This is consistent with how extra capacity is added in MoE. Similarly Li et al. (2024) present an ablation that shows that adding FFN blocks is better than adding attention blocks. Note, this is in contrast with what is observed by Hu et al. (2022) with the LoRA method, for which the best choice to adapt to different downstream tasks with few parameters is to update the MHA matrix, in particular they choose to update the keys and values matrices. In our preliminary experiments, we re-evaluate LoRA in our context of knowledge addition, and observe that LoRA is significantly more effective when updating the FFN blocks than the MHA blocks. We hypothesize that this apparent discrepancy is due to the fact that task adaptation is different from the question of adding a large body of knowledge to a pretrained network.
Mixed distribution training.
As discussed in the related work, one solution to limit catastrophic forgetting is to keep training with a proportion of the original data, if such data is available. This strategy is typically employed with finetuning but is also effective with all baselines. Li et al. (2024) report that catastrophic forgetting almost disappear when the volume of original language data is five times greater than the new one, yet they also point out that mixed data training is likely to slow down the learning of the new language.
In our work, we assume that we have access to a “similar distribution” at posttraining time when we learn the adapters. We utilize a proportion $p$ of data aimed at being similar to the original distribution (albeit not identical), which in our case is data in English. We also consider a case in which we do not know what kind of data was used at training time, like in the case of the Gemma et al. (2024) model. In our ablation section 4.4, we evaluate the impact of $p$ on learning and forgetting. As a particular case, we also analyze the case $p=0$, where no pretraining data is available or used.
Adapter gating & local loss.
Recent transformer models commonly adopt a gated activation with a SiLU non-linearity (Elfwing et al., 2018). The gating weights implement a linear operator $W_{\mathrm{g}}$. When training a vanilla adapter block for a new language, the singular values of this operator are significantly more skewed than in regular transformer block from the pretrained backbone. Figure 3 shows this behavior, which is detrimental to the learning process. Our interpretation is that all projections instantiated by this operator tend to be colinear and therefore to behave similarly when combined with the SiLU non-linearity, because they implicitly operate as a old/new data switch. The redundancy in the gating matrix hinders the capacity of the model.
To address the problem, we add a block gate as in MoE transformers (Shazeer et al., 2017) and two recent works on model extension with gated MoE (Zhong et al., 2024; Li et al., 2024). In our case, we use a single gate whose role is to operate as a selector, either explicit or implicit, coupled with a local loss applied at the block level. This gating is combined with two distinct strategies:
  • Explicit selector with sigmoidal activation. This solution is inspired by MoE. In this case, the block gate is associated with a sigmoid activation. It is therefore trained as a local binary classifier whose objective is to distinguish between the old and the new data distributions.
  • Implicit selector: ELU activation with sparsity constraint. This strategy, which is the one that we refer to as “neutral residues” unless specified otherwise, does not make an explicit classification. Instead, we let the ELU activation adapts itself the strength of its response for the block. We also apply a loss $\ell_{1}$ on the output of the adapter governed by a hyperparameter $\alpha$, which favors the adapter to return a residues of 0s. It is only employed when feeding the extended model with data of the original distribution.
As one can observe in Figure 3, the gating operator has a significant effect on the singular values of the gating matrix: the distribution of the singular values is less skewed than the original backbone. The implicit selector is more effective than the local classifier in that respect. As we will see in Section 4, it is the method that offers the best compromise between not forgetting and learning.
Figure 3: Spectral analysis of the gating matrix, normalized by the largest singular value.

### v3

@@ Abstract
We address the problem of extending a pretrained large language model to a new domain that was not seen during training. Standard techniques, such as finetuning or low-rank adaptation (LoRA) are successful at domain adaptation, but do not formally add capacity to the model. This often leads to a trade-off, between performing well on the new domain vs. degrading performance on the original domain. Here, we revisit and improve adapters to extend LLMs from three angles: data, architecture and training procedure, which are advantageously considered jointly. The resulting method, called neutral residues, modifies adapters in a way that leads each new residual block to output near-zeros on the original domain. This solution leads to strong results when adapting a state-of-the-art model originally trained on English to a new language. Neutral residues significantly outperform competing approaches such as finetuning, LoRA or vanilla adapters in terms of the trade-off between learning the new language and not forgetting English.
Machine Learning, ICML
Figure 1: Learning or Forgetting? Fine-tuning a model reduces the model performance on the original task. Potential trade-offs are governed by the learning rate, with extreme cases being (top-left) the original backbone and (bottom-right) finetuning on the new data with a large learning rate. LoRA or vanilla adapters mitigate catastrophic forgetting but only to some extent. Our method significantly improves this compromise. We report detailed results in Table 7 of Section 4.
@@ 1 Introduction
The dominating strategy for producing foundation models involves training from scratch on a large collection of data, typically trillions of tokens, covering multiple domains. Training such models is extremely costly in terms of computing resources. In view of the explosion of the number of gigantic models produced in the last years (the HuggingFace model repository claims to host about one million models), this training paradigm raises the problem of the economical and ecological sustainability of the model production pipelines. As an example, the cost of training the Llama 3 model is estimated be in the order of hundreds of millions dollars, mostly due to the training on 16,000 H100 GPUs for months, [^1] https://ai.meta.com/blog/meta-llama-3/ not counting the hidden costs of exploration.
@@
In contrast, the adapters introduced by Rebuffi et al. (2017) extend a pretrained model by adding new parameters. When transferring to a new domain or task, only these new weights are trained while those of the original backbone are frozen. This extension strategy, which was initially introduced in computer vision with convolutional neural networks, was subsequently adapted to textual transformers by Houlsby et al. (2019). Adapters benefit from the initial pretrained model and hence exhibit competitive transfer performance. They enjoy several additional properties. First, only a subset of weights is trained. Most importantly, they formally add capacity (unlike linear additive adapters like LoRA) and hence do not suffer from the aforementioned limitations. However, adapters still suffer significant forgetting in their current form and are therefore not sufficient for our problem.
In this paper, we build on adapters and, accordingly, increase the capacity of the model by adding feed-forward blocks, while offering a modular design (Pfeiffer et al., 2020). In particular, we consider the use-case where we add a new language to a pretrained model. We measure how the extended model performs on the training criterion, namely perplexity, but also on downstream tasks such as question answering.
We analyze how to optimize the compromise between learning a new language, and not forgetting the initial knowledge of the pretrained model. We study the impact of (1) training data, (2) training strategy, such as losses specifically intended at reducing the forgetting, and (3) architecture. This leads us to identify several important factors for successfully extending a model with adapters:
  • Data: When learning adapters, a simple way to reduce the forgetting is to keep training with a small fraction of data similar to the original distribution.
  • Architecture: An adapter gating mechanism is a way to ensure that the network distinguishes when it should operate as the original neural network or when the new blocks are desirable to process the input data.
  • Training: The gating mechanism is advantageously informed in a supervised manner. We proposed two strategies in that respect, both introducing a local loss: The first one is inspired by mixture of experts (MoE), which involves an explicit domain classifier at each block. The second one involves a sparsity loss whose objective is to ensure that the residual connections output near-zero values when the input follows the pretraining distribution.
@@
Overall, we show that these multiple ingredients are advantageously intertwined to obtain the best trade-off between learning the new distribution well and not forgetting the original capabilities. This leads to our proposed method, called neutral residues. We then apply neutral residues to improve the performance of the Gemma LLM (Gemma et al., 2024) on languages such as Danish, Hungarian or Slovak, without degrading the performance on English.
@@ 2 Related Work
Finetuning is however not sufficient when a large body of knowledge must be incorporated to the network. In this section we review a body of literature related to this problem and to our work. This includes solutions that add new knowledge into a pretrained network, as well as methods that incorporate gating mechanisms such as MoE.
@@
Enhancing language model capabilities using module based training has gained interest in recent years due to the high cost of full finetuning. Those methods are known as parameter efficient finetuning methods (PEFT; Houlsby et al. (2019); Hu et al. (2022); Li & Liang (2021)). Unlike full finetuning, they only require a limited amount of memory. Houlsby et al. (2019) proposed to insert small bottleneck layers, known as Adapters, within each transformer layer, allowing models to be finetuned by training only a small fraction of the parameters: The original model is frozen. There is a large body a literature on adapters, see the overview by Pfeiffer et al. (2023). LoRA (Hu et al., 2022) adds trainable low-rank matrices to transformer layers, while freezing the original weights.
@@
Several studies have investigated more advanced methods to address forgetting in continual learning (Biderman et al., 2024; Wu et al., 2024; Li et al., 2024; Zhong et al., 2024; Riemer et al., 2019; Li & Hoiem, 2017). Biderman et al. (2024) show that LoRA helps reducing forgetting but is less efficient at learning than full finetuning, especially for distributions far from the LLM pretraining distribution. Wu et al. (2024) proposed a continual learning method for LLMs, which amounts to adding new intermediate transformer blocks in an interleaved manner. This enables the injection of new knowledge while preserving the initial capabilities.
Mixture of Experts.
@@
MoEs have a long history in neural networks (Yuksel et al., 2012), for instance with set of classifiers where each classifier was specialized for a different task or domain. Residual architectures (He et al., 2016) and in particular transformers (Vaswani et al., 2017) have renewed the interest in this strategy. In this context, they amount to replacing the standard feed-forward networks (FFN) of transformer blocks by a collection of FFNs, referred to as experts. A gating mechanism selectively activates a subset of relevant experts for a given input. This technique has been widely used to reduce the computational cost of inference in large language models (Jiang et al., 2024; Xue et al., 2024) with a focus on the design of an expert-balancing loss to promote equitable token distribution across experts (Shazeer et al., 2017; Pfeiffer et al., 2023).
@@
Recent works have explored utilizing MoE’s gating mechanism to mitigate catastrophic forgetting during the continual learning phase of language models. LoRAMoE (Dou et al., 2024) employs LoRAs as experts, integrating them with a gating mechanism to improve performance on specific downstream tasks through Supervised finetuning (SFT). LoRAMoE mitigates catastrophic forgetting in the case of SFT, but to our knowledge its effectiveness has not been studied in a broader context. In contrast, two recent parallel concurrent works are closer to our proposal: Zhong et al. (2024) augment Mixture-of-Experts LLMs with new modalities by addition of new experts to a pretrained mixture of experts. Similarly, Li et al. (2024) add an MoE layer in parallel to the FFN of the transformer block to learn multilingual capabilities and mitigate catastrophic forgetting during Continual Learning. Our solution shares some similarities with this work by employing mixed data training. Yet Li et al. (2024) do not integrate an explicit sparsity criterion, and need to weight differently the blocks of the pretrained model with those of the gated adapters. They also consider multiple elements in the mixture and apply a softmax for the rooting mechanism. Since at least one expert is activated in their mixture, this selection mechanism also prevents the capacity of the model to produce 0-valued outputs. Our solution of neutral residues with local loss is more effective than simply gating adapters, as shown in our experimental Section 4.
@@ 3 Neutral Residues with Gated Adapters
The first adapters were initially introduced by Rebuffi et al. (2017), in the context of convolutional residual networks. Their motivation was to adapt a backbone to multiple tasks by finetuning on a limited amount of data, rather than adding a large body of knowledge. Therefore they considered a limited number of trainable parameters, typically less than 10% additional weights compared to the original backbone. In the context of language models, Houlsby et al. (2019) add residual feed-forward block sequentially after each multi-head attention and the original FFN. In our case and as shown in Figure 2, we add more capacity (typically 20%) and adopt parallel adapters as we do not observe a noticeable difference compared to serial adapters. This concurs with observations by Touvron et al. (2022) that blocks can be parallelized pairwise if vision transformers are large enough.
@@
Figure 2: Architectural design adopted for neutral residues: (left) we add the adapter block in parallel to the FFN ; (right) our gated adapter block includes a block gate with a ReLU non-linearity (see Section 4.4 for a comparison with a sigmoid-based gating). At training time, the weights of the blue blocks are frozen, while the trainable parameters are represented by red blocks. The output of the adapter is trained with a local loss such that the output is sparse if the input follows the pretraining distribution. We note that the $\sigma$ activation function of the adapter block is the same as the one used in the backbone.
Mixed distribution training.
As discussed in the related work, a solution to limit catastrophic forgetting is to keep training with a small amount of the original data, if such data is available. This strategy is typically employed with finetuning but is effective with all baselines. Li et al. (2024) report that catastrophic forgetting almost disappears when the volume of data in original language is five times greater than in the new one. Yet they also point out that mixed data training can slow down the learning of the new language.
Unfortunately, in many cases, the original training data is not available when extending a model. Thus, in the following, we do not assume that we have access to the original data distribution, but only to data that is somewhat related. For example, when using a model that was pretrained mostly on English Common Crawl, we use data from English Wikipedia to capture the original distribution.
We also consider the more realistic case of using a public model, for which we do not have details about the pretraining data distribution, such as the Gemma model (Gemma et al., 2024). We study the impact of the proportion $p$ of data related to the original domain. As a particuler case, we analyze the case $p=0$, where no data related to the pretraining is available.
Local loss & adapter gating.
One way to prevent forgetting the pretraining distribution is to constrain the output of the model so that it does not change on the original data. This can be achieved by forcing the outputs of the adapters to be equal to zero on the original data, and only non-zero on data from the new domain. To this end, we propose to regularize the output of the adapter modules with an $\ell_{1}$-norm on the data related to the original distribution. We argue that regularizing the output of adapters is easier than continual learning on the original distribution. Hence, using an imperfect dataset representing the original distribution should have a smaller impact on our method than finetuning.
Figure 3: Spectral analysis of the weight gating matrix of GLU blocks, normalized by the largest singular value.
During early experiments, we observed that regularizing the output of adapters helped mitigate forgetting, but it also prevented learning the new domain. As most recent transformer LLMs, the models used in our experiments use gated linear units (GLU) in feed-forward blocks (Dauphin et al., 2017; Shazeer, 2020). The output of the block is obtained with $\text{FFN}(x)~=~\mathbf{W}_{\mathrm{o}}\left(\sigma\left(\mathbf{W}_{\mathrm{g}}x\right)\odot\mathbf{W}_{\mathrm{i}}x\right),$ where $\odot$ is the pointwise multiplication and $\sigma$ is a non-linearity such as SiLU (Elfwing et al., 2018). We observed that the singular values of the gating weights $\mathbf{W}_{\mathrm{g}}$ of adapters are significantly more skewed than in regular transformer block from the pretrained backbone. Figure 3 shows this behavior, which is detrimental to the learning process. Our interpretation is that all the projections instantiated by this operator tend to be colinear, and therefore to behave similarly when combined with the non-linearity, because they implicitly operate as classifier between the old and new distributions.
To address this problem, we add a block gate over the full adapter, in a similar fashion as the gating mechanism of MoE transformers (Shazeer et al., 2017), and their application to model extension (Zhong et al., 2024; Li et al., 2024). In our case, we add a gate only on the additional adapter block, and not on the original block. We optimize the loss $\ell_{\text{train}}=\ell_{\text{LM}}+\alpha\ell_{\text{local}}$, where $\ell_{\text{LM}}$ is the language modeling objective and $\ell_{\text{local}}$ is a loss to train the gating to discriminate the original vs. the new distribution. We consider two gating activations and corresponding local losses:
  • ReLU activation with sparsity constraint. This strategy, which is the one that we refer to as “neutral residues” unless specified otherwise, does not make an explicit classification. Instead, we let the ReLU activation adapts itself the strength of its response for the block. We compute the average of the $\ell_{1}$-norm on the output of the adapters as the local loss. We normalize the loss by the dimension of the model. This loss is only used when training on the original distribution.
  • Sigmoid activation with cross entropy. This solution is inspired by MoE. In this case, the block gate is associated with the sigmoid activation. It is thus trained as a local binary classifier whose objective is to distinguish between the old and the new data distributions, using the cross-entropy loss averaged over the adapters. In contrast to the previous strategy, this loss is used on both the original and the new distributions.
In Figure 3, observe that the gating operator has a significant effect on the singular values of the gating matrix: the distribution of the singular values is less skewed than the original backbone. Moreover, the gating mechanism based on ReLU and $\ell_{1}$-norm is more effective than the one based on sigmoid and the cross-entropy loss. As we will see in Section 4, it is also the method that offers the best compromise between not forgetting and learning.
Low-variance initialization.
@@
FFN vs. multi-head attention (MHA).
In our early experiments carried out to add a new language, it is more effective to add extra weights in the form of FFN than in MHA blocks, for a given number of additional parameters. This is consistent with how extra capacity is added in MoE. Similarly Li et al. (2024) present an ablation that shows that adding FFN blocks is better than adding attention blocks. Note, this is in contrast with what is observed by Hu et al. (2022) with the LoRA method, for which the best choice to adapt to different downstream tasks with few parameters is to update the MHA matrix, in particular they choose to update the keys and values matrices. In our preliminary experiments, we re-evaluate LoRA in our context of knowledge addition, and observe that LoRA is more effective when updating the FFN blocks than the MHA blocks. We hypothesize that this apparent discrepancy is due to the fact that task adaptation is different from the task of extending a pretrained network on a large body of knowledge.